OpenStudy (anonymous):

Can you simplify 256/625?

OpenStudy (anonymous):

nope, that's the most simplified that it can be

OpenStudy (anonymous):

No simplification possible

OpenStudy (anonymous):

Alright cool thanks guys!

OpenStudy (tkhunny):

"Simplified" is not a particular clear concept. \(\dfrac{256}{625} = \dfrac{2^{8}}{5^{4}} = \left(\dfrac{4}{5}\right)^{4}\). That looks "simplified" to me.